# CST Anti-UAV: A Thermal Infrared Benchmark for Tiny UAV Tracking in Complex Scenes

### v1

@@ 4.1 Experimental Settings
Evaluation metrics. We use three common evaluation metrics, namely precision plot, success plot, and state accuracy, to assess tracker performance through One-Pass Evaluation (OPE). The calculation methods for precision plots and success plots are the same as those in [50]. State accuracy consistent with the previous method [29], represents the Intersection over Union (IoU) between the predicted bounding box and its corresponding ground-truth, while denotes the visibility flags of the ground-truth annotations. The tracker’s predicted is used to measure the state accuracy. The average state accuracy of all video sequences is the mSA.
Baseline tracker. We conducted comprehensive evaluations of 20 single object trackers on both CST Anti-UAV and Anti-UAV410 [26] datasets. The transformer trackers include Stark [57], Stark-ST [57], OSTrack [59], AiATrack [23], MixFormer [13], MixFormerV2 [14], DropTrack [49], ARTrack [48], SeqTrack [10], GRM [24], ROMTrack [6], HIPTrack [5], AQATrack [55] and ARTrackV2 [1]. The other deep-learning trackers contain GlobalTrack [27], KYS [4], PrDiMP [17], ToMP [38], and Refocus-TIR [31]. For the Siamese trackers, we have only selected SiamDT [26], which is designed for TIR UAV tracking tasks. We conducted all experiments on a server equipped with 8 NVIDIA A100-PCIE-40GB GPUs and used the implementations, pre-trained models, and default configurations provided by the official code to ensure a fair comparison among the trackers.
@@ 4.2 Overall Performance
Table 3 presents the results on both Anti-UAV410 and CST Anti-UAV datasets, including state accuracy, and the Area Under the Curve (AUC) of success and precision plots. Notably, trackers that achieve superior performance on Anti-UAV410 exhibit significant degradation when evaluated on the proposed CST Anti-UAV dataset. The poor performance on CST Anti-UAV is not only due to complex backgrounds and tiny objects in static images but also because of object disappearance and reappearance across the temporal domain. Based on the evaluation, SiamDT [26] and GlobalTrack [27] secured the top two positions with an absolute advantage on both datasets. This indicates that trackers based on long-term tracking have the potential to achieve better performance, as they search in the whole image, allowing them to obtain the disappeared object’s reappearing position.
In addition to experiments between different trackers, we further explored the impact of various tracker versions and different search region sizes on their performance. Stark-ST, which adds a dynamically changing template to capture temporal variations in targets, achieved an improvement in performance. This demonstrates that effectively utilizing temporal information can help address changes in object size and appearance. OSTrack-384 [59] and ARTrack-384 [48] show notable performance improvements compared to OSTrack-256 and ARTrack-256 on Anti-UAV410 but have limited impact on CST Anti-UAV. We attribute this to the tiny size of the targets, which occupy fewer pixels in the image, making it harder for the network to effectively leverage its semantic representation capabilities.
@@
| GlobalTrack [27] | 30.96 | 35.92(+16.02%) |
| OSTrack-256 [59] | 21.81 | 26.34(+20.77%) |
| OSTrack-384 [59] | 21.29 | 25.76(+21.00%) |
| SeqTrack [10] | 20.36 | 26.53(+30.30%) |
| GRM [24] | 21.64 | 26.96(+24.58%) |
| ROMTrack [6] | 21.13 | 27.43(+29.82%) |
| AQATrack [55] | 23.12 | 27.60(+19.38%) |
| ARTrackV2 [1] | 19.14 | 22.14(+15.67%) |
| SiamDT [26] | 30.95 | 35.84(+15.80%) |
| Refocus-TIR [31] | 24.90 | 26.51(+6.47%) |
@@ 4.4 Coarse-to-fine Attribute Evaluation
Concretely, GlobalTrack [27] and SiamDT [26] still perform well across various attributes and object scales except for tiny objects, likely due to their global search modules. While this ensures strong target re-localization capability after the disappearance, the expanded search regions introduce more background noise, leading to degraded performance on tiny-sized objects. KYS [4] uses dense localized state vectors and a prediction module to propagate scene information across frames, enabling strong performance in tracking out-of-view attributes and tiny objects. PrDiMP [17] combines probabilistic regression and DiMP [3] to provide a probabilistic distribution prediction for the target’s location, which allows precise localization after occlusions and adapts to the various object scales. PrDiMP shows strong performance on challenges such as occlusions, out-of-view, and complex dynamic backgrounds, and achieves state-of-the-art results for tiny-sized objects.

### v2

@@ 4.1 Experimental Settings
Evaluation metrics. We use three common evaluation metrics, namely precision plot, success plot, and state accuracy, to assess tracker performance through One-Pass Evaluation (OPE). The calculation methods for precision plots and success plots are the same as those in [otb50]. State accuracy consistent with the previous method [antiuav], represents the Intersection over Union (IoU) between the predicted bounding box and its corresponding ground-truth, while denotes the visibility flags of the ground-truth annotations. The tracker’s predicted is used to measure the state accuracy. The average state accuracy of all video sequences is the mSA.
Baseline tracker. We conducted comprehensive evaluations of 20 single object trackers on both CST Anti-UAV and Anti-UAV410 [antiuav410] datasets. The transformer trackers include Stark [stark], Stark-ST [stark], OSTrack [ostrack], AiATrack [aiatrack], MixFormer [mixformer], MixFormerV2 [mixformerv2], DropTrack [droptrack], ARTrack [artrack], SeqTrack [seqtrack], GRM [grm], ROMTrack [romtrack], HIPTrack [hiptrack], AQATrack [aqatrack] and ARTrackV2 [artrackv2]. The other deep-learning trackers contain GlobalTrack [globaltrack], KYS [kys], PrDiMP [prdimp], ToMP [tomp], and Refocus-TIR [refocus]. For the Siamese trackers, we have only selected SiamDT [antiuav410], which is designed for TIR UAV tracking tasks. We conducted all experiments on a server equipped with 8 NVIDIA A100-PCIE-40GB GPUs and used the implementations, pre-trained models, and default configurations provided by the official code to ensure a fair comparison among the trackers.
@@ 4.2 Overall Performance
Table 3 presents the results on both Anti-UAV410 and CST Anti-UAV datasets, including state accuracy, and the Area Under the Curve (AUC) of success and precision plots. Notably, trackers that achieve superior performance on Anti-UAV410 exhibit significant degradation when evaluated on the proposed CST Anti-UAV dataset. The poor performance on CST Anti-UAV is not only due to complex backgrounds and tiny objects in static images but also because of object disappearance and reappearance across the temporal domain. Based on the evaluation, SiamDT [antiuav410] and GlobalTrack [globaltrack] secured the top two positions with an absolute advantage on both datasets. This indicates that trackers based on long-term tracking have the potential to achieve better performance, as they search in the whole image, allowing them to obtain the disappeared object’s reappearing position.
In addition to experiments between different trackers, we further explored the impact of various tracker versions and different search region sizes on their performance. Stark-ST, which adds a dynamically changing template to capture temporal variations in targets, achieved an improvement in performance. This demonstrates that effectively utilizing temporal information can help address changes in object size and appearance. OSTrack-384 [ostrack] and ARTrack-384 [artrack] show notable performance improvements compared to OSTrack-256 and ARTrack-256 on Anti-UAV410 but have limited impact on CST Anti-UAV. We attribute this to the tiny size of the targets, which occupy fewer pixels in the image, making it harder for the network to effectively leverage its semantic representation capabilities.
@@
| GlobalTrack [globaltrack] | 30.96 | 35.92(+16.02%) |
| OSTrack-256 [ostrack] | 21.81 | 26.34(+20.77%) |
| OSTrack-384 [ostrack] | 21.29 | 25.76(+21.00%) |
| SeqTrack [seqtrack] | 20.36 | 26.53(+30.30%) |
| GRM [grm] | 21.64 | 26.96(+24.58%) |
| ROMTrack [romtrack] | 21.13 | 27.43(+29.82%) |
| AQATrack [aqatrack] | 23.12 | 27.60(+19.38%) |
| ARTrackV2 [artrackv2] | 19.14 | 22.14(+15.67%) |
| SiamDT [antiuav410] | 30.95 | 35.84(+15.80%) |
| Refocus-TIR [refocus] | 24.90 | 26.51(+6.47%) |
@@ 4.4 Coarse-to-fine Attribute Evaluation
Concretely, GlobalTrack [globaltrack] and SiamDT [antiuav410] still perform well across various attributes and object scales except for tiny objects, likely due to their global search modules. While this ensures strong target re-localization capability after the disappearance, the expanded search regions introduce more background noise, leading to degraded performance on tiny-sized objects. KYS [kys] uses dense localized state vectors and a prediction module to propagate scene information across frames, enabling strong performance in tracking out-of-view attributes and tiny objects. PrDiMP [prdimp] combines probabilistic regression and DiMP [dimp] to provide a probabilistic distribution prediction for the target’s location, which allows precise localization after occlusions and adapts to the various object scales. PrDiMP shows strong performance on challenges such as occlusions, out-of-view, and complex dynamic backgrounds, and achieves state-of-the-art results for tiny-sized objects.